MAKEMPX(1) MAKEMPX(1) NAME makempx−typesetlabelsinMetaPostpictureswithTeXorTroff SYNOPSIS makempx[--tex|--troff]mpfilempxfile DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Ifmpxfileisolderthanmpfile,translatethelabelsfromtheMetaPostinputfilempfiletolow-level commandsinMPXFILE,byrunningmpto-tex,tex,anddvitompbydefault;or,if-troffis specified,mpto-troff,eqn-d\$\$|troff-Tpost,anddmp. Thecurrentdirectoryisusedforwritingtemporaryfiles.Errorsareleftinmpxerr.{tex,log,dvi}. Ifthefilenamedin$MPTEXPRE(mptexpre.texbydefault)exists,itisprependedtotheoutput intexmode. makempxistypicallycalledbympost,notfromthecommandline. OPTIONS makempxacceptsthefollowingoptions: --helpPrinthelpmessageandexit. --texAssumeinputfilecontainsTEXcommands. --troffAssumeinputfilecontainsTroffcommands. --version Printversioninformationandexit. SEEALSO mpost(1),mpto(1). Web2C7.32x 4January1998 1